Your company has earnings per share of $4. It has 1 million shares outstanding, each of which has a price of $40. You are thijust need part d please show equations
Answer #1

Part d)

i) Value of the acquirer pre merger

= market price per share * no. of shares outstanding of acquirer

= $ 40 * 1 million = $ 40 million

ii) Value of the target pre merger

= market price per share * no. of shares outstanding of target

= $ 25 * 1 million = $ 25 million

iii) total equity value after merger

= Market price per share * (no. of shares outstanding of acquirer + no. of shares issued to Target)

= $40 * (1 million + { 1 million * (2/4)})

= $40 * 1.5 million shares

= $ 60 million

1a)Stock price after merger,

Earning per share after merger

= Total earnings / after merger no. of shares outstanding

= ( {$4 * 1 million shares of acquirer} + {$2 * 1 million shares of target}) / ( 1 million shares of acquirer + { 1 million shares of target * exchange ratio 2/4}

= $ 6 million / 1.5 million after merger shares

= $ 4

P/E ratio after merger = stock price after merger / earnings per share after merger

Therefore stock price after merger = earning per share after merger * PE ratio after merger

= $4 * 10

= $ 40

1b) P/E ratio after merger = Earnings per share / stock price after merger

= $40 / $ 4

= 10 times

1c) P/E ratio of acquirer pre merger

= $40 / $ 4

= 10 times

1d) P/E ratio of target pre merger

= $ 25 / $2

= 12.50 times

There is no change in the P/E ratio of pre merger and after merger since the exchange ratio of merger is calculated based on the earnings per share